OpenStudy (anonymous):

What is the y-value of cos(x) when x = -360°? A. -1 B. 0 C. one-half D. 1

OpenStudy (anonymous):

\[\huge \cos(- \theta) = \cos(\theta)\]

OpenStudy (anonymous):

So, \[\cos(-360) = \cos(360)\]

OpenStudy (anonymous):

and: \[\cos(360) = \cos(0) = 1\]

OpenStudy (anonymous):

so the answer is 1

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):